# Probing black hole entropy via entanglement

### v1

@@ 1 Introduction
Figure 3: The shaded square regions on both sides denote equivalent regions of interest. The left-hand side of the figure shows the Penrose diagram of a $D$-dimensional extremal RN black hole. The strip enclosed by dashed purple lines represents its near-horizon region, which takes the form AdS2×$\times S^{D-2}$. Each point along the bold black line corresponds to the same area of the event horizon. The Bekenstein–Hawking entropy associated with the black hole is entirely determined by the AdS2 geometry, captured by the blue dot labeled $\mathrm{Area}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma}\right)$. In the entropy, the higher-dimensional spherical part $S^{D-2}$ of the spacetime metric is absorbed into the $D$-dimensional Newton’s constant $G_{N}^{\left(D\right)}$, effectively reducing it to a two-dimensional Newton’s constant, denoted $G_{N}^{\left(2\right)}$. The right-hand side of the figure illustrates the emergent AdS2 spacetime derived from the entanglement entropy of a TFD state. In this setup, the RT prescription yields the area associated with the red dot $\mathrm{Area}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma\left(A:B\right)}\right)$. Since the two purple strip regions represent the same near-horizon geometry, the areas can be naturally identified: $\mathrm{Area}\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\Sigma}\right)=\mathrm{Area}\left({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\gamma\left(A:B\right)}\right)$. This observation implies that the $D$-dimensional Bekenstein–Hawking entropy of the extremal black hole can be recovered from the entanglement entropy of CQM1.
@@ 2.1 BTZ black hole
[…]
@@
Remarkably, the entropy of system 1 is precisely the entanglement entropy between the two CFT copies, which can be given by:
@@ 2.2 $D1$-$D5$ black hole of type IIB string theory
A quick example is the $D1$-$D5$ black hole, which arises as a solution of ten-dimensional type IIB string theory with $Q_{1}$ D1111-branes and $Q_{5}$ D5555-branes wrapped on a compact internal space $T^{4}\times S^{1}$. Here, the four-torus $T^{4}$ has volume $\left(2\pi\right)^{4}v\alpha^{\prime}$, and the circle $S^{1}$ has radius $R$ [18]. The Kaluza-Klein momentum along $S^{1}$ is quantized and labeled by $N$. The near-horizon limit is obtained by taking the string length squared $\alpha^{\prime}\rightarrow 0$, leading to the following geometry:
@@
Moreover, it is worth noting that the near horizon limit of near extremal BTZ is AdS2×$\times S^{1}$. This implies that the corresponding TFD state in AdS2/CFT1 can also be used to compute the entanglement entropy. In this case, one starts with a rotating BTZ black hole, which modifies the CFT defined on the asymptotic boundary. The TFD state must then be replaced by [7]:
@@
[…]
@@ 3.1.1 RN black hole at large $D$
Let us begin with the first step circled-1, where we aim to demonstrate that the large $D$ limit of the RN black hole reduces to the two-dimensional charged black hole solution in string theory. We first recall the $D$-dimensional RN black hole solution in Einstein–Maxwell theory [21]:
@@
Now, we consider step circled-2, which involves the action description corresponding to the black hole solution obtained in step circled-1. We begin with the $D$-dimensional Einstein–Maxwell theory, whose action is given by:
@@
[…]
@@ 3.1.2 Extremal limit of RN black hole at large $D$
In this subsection, we proceed to discuss steps circled-3 and circled-4, which involve taking the extremal limit of the large $D$ black hole and analyzing its associated action. Recall that the inner and outer horizons of a large $D$ black hole are determined by the condition:
@@
The field strength squared becomes: $F^{2}=-\frac{4\lambda^{2}q^{2}}{\bar{\mathrm{R}}^{2}}=-4\lambda^{2}$. We observe that the $\left(\tau,\rho\right)$ sector of the geometry is precisely AdS2, completing the near-horizon, extremal limit of the large $D$ charged black hole.
@@ 3.2 CQM1 entanglement entropy
which exactly reproduces the Bekenstein–Hawking entropy (3.63) of the large $D$ extremal RN black hole.
@@
where $\lambda=\frac{n}{2r_{0}}$. Using the classical solutions $\phi=\mathrm{const}.$, $A_{t}=\sqrt{\frac{1}{2}}\frac{n}{r_{0}}\left(\rho-\frac{1}{2}\right)d\tau,$ and the field strength $F^{2}=-\frac{4\lambda^{2}q^{2}}{\bar{\mathrm{R}}^{2}}=-4\lambda^{2}$, one introduces the replica geometry via the curvature singularity $\mathbb{R}=4\pi\left(1-N\right)\delta^{2}\left(x\right)$. This leads to the action: $I_{\mathrm{string}}=\frac{N-1}{4G_{N}^{\left(2\right)}}$. The entanglement entropy is then computed by
@@
Moreover, the same result is supported by recent developments in the nearly-AdS2/SYK correspondence [24].

### v2

@@ 1 Introduction
[…]
@@ 2.1 BTZ black hole
$\left|TFD\right\rangle=\frac{1}{\sqrt{Z\left(\beta\right)}}\underset{n}{\sum}e^{-\beta E_{n}/2}\left|E_{n}\right\rangle_{1}\otimes\left|E_{n}\right\rangle_{2},\qquad Z\left(\beta\right)=\underset{n}{\sum}e^{-\beta E_{n}},$ (2.5)
@@
The entropy of system 1 is precisely the entanglement entropy between the two CFT copies, which can be given by:
@@ 2.2 $D1$-$D5$ black hole of type IIB string theory
A quick example is the $D1$-$D5$ black hole, which arises as a solution of ten-dimensional type IIB string theory with $Q_{1}$ D11-branes and $Q_{5}$ D55-branes wrapped on a compact internal space $T^{4}\times S^{1}$. Here, the four-torus $T^{4}$ has volume $\left(2\pi\right)^{4}v\alpha^{\prime}$, and the circle $S^{1}$ has radius $R$ [18]. The Kaluza-Klein momentum along $S^{1}$ is quantized and labeled by $N$. The near-horizon limit is obtained by taking the string length squared $\alpha^{\prime}\rightarrow 0$, leading to the following geometry:
@@
[…]
@@
$\left|TFD\right\rangle=\frac{1}{\sqrt{d\left(N\right)}}\underset{n}{\sum}\underset{k=1}{\overset{d\left(N\right)}{\sum}}e^{-\frac{\beta E_{n}}{2}}\left|k_{L},n_{R}\right\rangle_{\mathrm{CQM}1}\otimes\left|k_{L},n_{R}\right\rangle_{\mathrm{CQM}2},$ (2.23)
@@ 3.1.1 RN black hole at large $D$
Let us begin with the first step , where we aim to demonstrate that the large $D$ limit of the RN black hole reduces to the two-dimensional charged black hole solution in string theory. We first recall the $D$-dimensional RN black hole solution in Einstein–Maxwell theory [21]:
@@
Now, we consider step , which involves the action description corresponding to the black hole solution obtained in step . We begin with the $D$-dimensional Einstein–Maxwell theory, whose action is given by:
@@
where $\mathbb{R}$ is the Ricci scalar associated with the 2DD metric, and the volume of the unit $\left(n+1\right)$-sphere is given by $\Omega_{n+1}=2\pi^{\frac{n+2}{2}}/\Gamma\left(\frac{n+2}{2}\right)$. Taking the large $n$ limit ($n\rightarrow\infty$), the action reduces to the two-dimensional string effective action:
@@ 3.1.2 Extremal limit of RN black hole at large $D$
In this subsection, we proceed to discuss steps and , which involve taking the extremal limit of the large $D$ black hole and analyzing its associated action. Recall that the inner and outer horizons of a large $D$ black hole are determined by the condition:
@@
The field strength squared becomes: $F^{2}=-\frac{4\lambda^{2}q^{2}}{\bar{\mathrm{R}}^{2}}=-4\lambda^{2}$. We observe that the complete metric reduces to $\mathrm{AdS_{2}\times S^{D-2}}$, where the $\left(\tau,\rho\right)$ sector is precisely AdS2. The factor $r_{0}^{2}$ multiplying $d\Omega_{D-2}^{2}$ indicates that this geometry describes the near-horizon region around $r\rightarrow r_{0}$, with the transverse sphere remaining of fixed radius. It is important to emphasize that this AdS2 near-horizon geometry arises only in the extremal limit. For non-extremal black holes, the near-horizon geometry is approximately $\mathrm{Rindler\times S^{D-2}}$, rather than $\mathrm{AdS_{2}\times S^{D-2}}$. Thus, the appearance of the AdS2 throat is a direct consequence of extremality. Moreover, the spherical part $S^{D-2}$ does not affect the conformal boundary structure of the near-horizon geometry. Near the AdS2 boundary, the conformal factor of the two-dimensional metric diverges, while the sphere maintains finite size $r_{0}^{2}$. After dimensional reduction, the entire contribution of the sphere is absorbed into the effective two-dimensional Newton constant $G_{N}^{\left(2\right)}$. Consequently, the entropy is controlled by $1/G_{N}^{\left(2\right)}$, and the original horizon “area” is encoded in the two-dimensional theory rather than appearing as an explicit geometric area. In the reduced AdS2 description, the event horizon corresponds to a point on a constant-time slice. In this sense, the AdS2 Penrose diagram can be viewed as a zoomed-in description of the near-horizon region of the Reissner–Nordström geometry, as illustrated in figure (3). Since AdS2 possesses two asymptotic boundaries, the near-horizon limit isolates a two-boundary structure from the full RN Penrose diagram. In this limit, the black hole horizon of the RN geometry is mapped to the Poincaré horizon of AdS2.
@@ 3.1.3 Bekenstein-Hawking entropy
Therefore, the result follows form ollows from integrating out the transverse sphere in the dimensional reduction ansatz. The factor $\Omega_{n+1}r_{0}^{n+1}$ corresponds precisely to the horizon area contribution of the internal sphere, which becomes renormalized into the effective gravitational coupling of the two-dimensional theory.
@@ 3.2 CQM1 entanglement entropy
which exactly reproduces the Bekenstein–Hawking entropy (3.63) of the large $D$ extremal RN black hole. It is important to emphasize that extremality is crucial in our analysis. The emergence of an AdS2 factor in the near-horizon region is a universal property of extremal black holes. This structure ensures the decoupling of the near-horizon throat from the asymptotic region and leads to a two-boundary Penrose diagram, thereby making the AdS2/CQM1 framework applicable. Without extremality, the near-horizon geometry does not factorize into $\mathrm{AdS_{2}\times S^{D-2}}$. Instead, it is approximately $\mathrm{Rindler\times S^{D-2}}$, and the entropy contains non-universal thermal contributions associated with finite temperature excitations. In such cases, one does not expect the Bekenstein–Hawking entropy to be entirely captured by pure inter-boundary entanglement in a CQM1 description. Furthermore, in time-dependent or dynamical situations, the appropriate framework would be the covariant Hubeny–Rangamani–Takayanagi (HRT) prescription rather than the static RT surface. Whether the present identification between black hole entropy and inter-boundary entanglement entropy can be extended to the covariant case remains an interesting direction for future investigation.
@@
where $\lambda=\frac{n}{2r_{0}}$. Using the classical solutions $\phi=\phi_{0}=\mathrm{const}.$, $A_{t}=\sqrt{\frac{1}{2}}\frac{n}{r_{0}}\left(\rho-\frac{1}{2}\right)d\tau,$ and the field strength $F^{2}=-\frac{4\lambda^{2}q^{2}}{\bar{\mathrm{R}}^{2}}=-4\lambda^{2}$, one introduces the replica geometry via the curvature singularity $\mathbb{R}=4\pi\left(1-N\right)\delta^{2}\left(x\right)$. This leads to the action: $I_{\mathrm{string}}=\frac{N-1}{4G_{N}^{\left(2\right)}}$. The entanglement entropy is then computed by
@@
In this work, the large $D$ reduction of Einstein–Maxwell theory leads to the two-dimensional heterotic string effective action (3.67). In the extremal limit, the near-horizon solution reduces to a two-dimensional dilaton gravity theory equivalent to Jackiw–Teitelboim (JT) gravity without the topological term [15]. In standard JT gravity, the black hole entropy is determined by the value of the dilaton at the horizon. In our setup, after taking the extremal and large $D$ limits, the dilaton profile becomes constant $\phi_{0}$ in the near horizon region, so that $\exp\left(-2\phi_{0}\right)$ acts as an overall multiplicative factor in the action. This constant factor effectively renormalizes the two-dimensional Newton constant, yielding the effective coupling $G_{N}^{\left(2\right)}$. Consequently, the entropy reduces to $S=1/4G_{N}^{\left(2\right)}$, which matches the standard JT gravity expression when the horizon dilaton value is identified with the constant background. This result is consistent with the nearly-AdS2/SYK correspondence [24], where the entropy is dominated by the infrared AdS2 region and exhibits universal behavior governed by effective two-dimensional dilaton gravity. In particular, the large $D$ reduction isolates precisely this IR AdS2 sector, explaining why the entropy is fully captured by the two-dimensional theory. We emphasize that our analysis relies on this effective two-dimensional gravitational description (3.67) obtained from the large $D$ reduction. The consistency of such AdS2 dilaton gravity frameworks has been extensively validated in JT gravity and related models, providing a solid conceptual foundation for applying the entanglement-based interpretation in the present context.
Finally, we would like to place our results within the broader program of entanglement-based interpretations of black hole entropy. In our recent work [25], we identified a deeper mechanism underlying the statement that entanglement accounts for horizon degrees of freedom. Specifically, the entanglement entropy admits an effective two-dimensional description whose near-coincidence limit reduces to a string worldsheet theory propagating in a curved background. Requiring quantum consistency of this worldsheet theory leads to background field equations equivalent to Einstein gravity in AdS, and necessitates the presence of the spacetime metric, an antisymmetric Kalb–Ramond two-form field, and a dilaton.
In the standard holographic framework, entanglement entropy is computed geometrically via the RT surface and is therefore expressed in terms of the spacetime metric. However, within the equivalent string-theoretic description, the same quantity can be reformulated in terms of a divergenceless Kalb–Ramond two-form flux. In our construction, the associated conserved charge flow—sourced by ensembles of open strings—reproduces known bit-thread configurations and saturates the max-flow/min-cut bound. This establishes an explicit equivalence between entanglement entropy and open string charge flux.
Through open–closed string duality, this open string description is dual to a closed string sector that governs the bulk gravitational dynamics. In this sense, the Bekenstein–Hawking entropy, geometrically given by the area law, admits an equivalent interpretation in terms of closed string charge. The present work provides a concrete realization of this open closed correspondence in a holographic setting where the entropy can be computed in a fully controlled manner.
This perspective is conceptually aligned with earlier ideas of Susskind and Uglum [1], who proposed that black hole entropy arises from quantum degrees of freedom near the horizon. Here, that intuition is embedded within a precise holographic framework in which horizon entropy is explicitly identified with entanglement encoded in string-theoretic degrees of freedom.